# Explainable bank failure prediction models: Counterfactual explanations to reduce the failure risk

### v1

@@ 1 Introduction
Banking plays a significant role in the economic system, and any issues within this sector can have a more substantial adverse impact on the economic system than expected [1]. Even the failure of a single bank can trigger a chain reaction, potentially spreading swiftly and adversely affecting other banks [2]. Therefore, predicting problems within the banking system in advance has become an essential concern for the scientific community. One such effort is bank failure prediction models employed to forecast the likelihood of a bank’s failure using its financial indicators. Like in many other fields, the logistic regression model has been the baseline in bank failure prediction models for many years. This is due to its ease of creation and inherently interpretable structure, making it easily understandable by stakeholders. However, the non-linear complex relationships among variables in the models led to the adoption of more complex models as alternatives to logistic regression. Consequently, decision trees, support vector machines, and deep learning models started to be utilized in bank failure prediction models [1, 3, 4, 5, 6]. Despite exhibiting high prediction performance, these models were overshadowed by their lack of inherently interpretable structures. This is because laws such as GDPR [^1] Regulation (EU) 2016/679 of the European Parliament and of the Council of 27 April 2016 on the protection of natural persons concerning the processing of personal data and on the free movement of such data, and repealing Directive 95/46/EC (General Data Protection Regulation) https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX%3A32016R0679 and the AI Act [^2] Proposal for a Regulation of the European Parliament and of the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX%3A52021PC0206 imposed sanctions and restrictions on the use of such models due to their unexplainable nature.
@@
In recent years, explainable artificial intelligence (XAI) tools have been employed to open the black boxes, in other words, to explain opaque models [7, 8]. These tools have rapidly developed to solve the uninterpretable model structures in finance. The preliminary work by [9] investigated the importance of the variables in the bank failure prediction models, and then [5] obtained the most important variables similarly. [10] explains their financial distress models for the companies to obtain the important variables that can be used to treat loan eligibility requirements. [11] explains the financial distress model to meet the needs of external users generated by local and global explanations. [12] explain the models that are used to predict the financial distress of the companies in China to ensure the reliability of model outputs. While there are studies on explaining black box models in finance, studies on bank failure prediction models have focused only on variable importance. Therefore, it is clear that there is a need for research on explainability in terms of understanding and accountability of such models. Not only have the explainable models, but it is also utilized to derive actionable insights from the models [13, 14]: counterfactual explanations (CE), which describe the changes in the variable values needed to flip the prediction in the intended direction, are used to be actionable. The CE is one of the popular and easy-to-understand XAI tools in many domains such as healthcare [15], education [16, 17], finance [18, 19, 20, 21]. It is an output of an optimization process and is generated with methods such as What-If [22], Nearest instance [23], and Multi-objective counterfactual explanation [24]. The number of CEs that can be generated is potentially limitless [25]; however, generating high-quality CEs can be very difficult [26] because they are quite rare in most cases [27]. The high-quality CEs should be plausible, proximity, sparse, and valid [27]. There are preliminary studies to explore the desired properties of CEs in terms of desiderata. [28] indicated that the quality counterfactuals should be sparse with two variable changes from a psychological point of view. [29] proposed using more plausible CEs instead of proximity CEs because the plausible CEs are more robust. However, no uniformly better CE generation method exists for all datasets from different domains [30]. Thus, it is necessary to evaluate quantitatively [31, 32] the generated CEs by different methods before deployment.
@@
Additionally, the challenges encountered during the model selection phase include poor out-of-time performance and the issue of imbalanced data. The diminished out-of-time performance of bank failure prediction models results in their swift obsolescence [33, 34]. Hence, besides training models with robust out-of-sample performance, ensuring high out-of-time performance for long-term horizon [35, 5] is also imperative. Conversely, the issue of imbalanced data surfaces when there’s a notable scarcity of data examples for failed banks, leading to misclassification by the bank failure prediction model, particularly for the minority class. From the data-centric AI perspective, various resampling methods such as oversampling, undersampling, and Synthetic Minority Over-sampling Technique [36] are employed to mitigate this. However, the efficacy of resampling methods has come under scrutiny in recent years due to their adverse effects on the model [37, 38, 39, 40, 41], sparking debates regarding their utility [42]. Thus, we lean towards using the cost-sensitive approach rather than the resampling methods to address the imbalanced data. Consequently, the impact of resampling techniques on models has been rigorously assessed on both out-of-sample and out-of-time datasets, compared with results obtained using cost-sensitive approaches.
@@
Support vector machines, neural networks, and discriminant analysis are frequently enlisted in bank failure prediction modeling. However, despite their established efficacy in tabular datasets, tree-based models have not garnered widespread adoption, in contrast to deep learning networks [4, 5, 43]. Hence, we offer a novel perspective by leveraging tree-based models, namely decision trees [44], random forests [45], and extra trees [46]. The rationale behind employing multiple models lies in their diverse predictive capabilities, yielding different outcomes with varying variances. This paper solves basic bank failure prediction problems such as imbalance, out-of-sample, and out-of-time performance for US banks. Then, it determines the most appropriate method to generate actionable explanations from black-box bank failure prediction models, considering the effects of the balancing strategies used to produce the highest quality counterfactual explanations. To our knowledge, this is the first paper that proposes using counterfactual explanations for extracting actionable insights from the bank failure prediction models. This provides valuable insights to reduce the failure risk of banks, in addition to understanding the model. Moreover, we solve the problems of imbalancedness and out-of-time performance while obtaining the desired counterfactuals. In summary, our contribution is: (1) to train an accurate bank failure prediction model for the US banks against addressing imbalanced data and long time horizon, (2) to obtain the best counterfactual generation method for the bank failure prediction model, and (3) to illustrate how to use the counterfactual explanations to avoid bank failure in practice. This paper is organized as follows. Section 2 provides an overview of the counterfactual explanations, and Section 3 provides information about the conducted experiments on counterfactual generation methods and presents the results. The last section summarises the findings in the paper and suggests further studies.

### v2

@@ 1 Introduction
Banking plays a significant role in the economic system, and any issues within this sector can have a more substantial adverse impact on the economic system than expected [1]. Even the failure of a single bank can trigger a chain reaction, potentially spreading swiftly and adversely affecting other banks [2]. Therefore, predicting problems within the banking system in advance has become an essential concern for the scientific community. One such effort is bank failure prediction models employed to forecast the likelihood of a bank’s failure using its financial indicators. Like in many other fields, the logistic regression model has been the baseline in bank failure prediction models for many years. This is due to its ease of creation and inherently interpretable structure, making it easily understandable by stakeholders. However, the non-linear complex relationships among variables in the models led to the adoption of more complicated models as alternatives to logistic regression. Consequently, decision trees, support vector machines, and deep learning models started to be utilized in bank failure prediction models [1, 3, 4, 5, 6]. Despite exhibiting high prediction performance, these models were overshadowed by their lack of inherently interpretable structures. This is because laws such as GDPR [^1] Regulation (EU) 2016/679 of the European Parliament and of the Council of 27 April 2016 on the protection of natural persons concerning the processing of personal data and on the free movement of such data, and repealing Directive 95/46/EC (General Data Protection Regulation) https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX%3A32016R0679 and the AI Act [^2] Proposal for a Regulation of the European Parliament and of the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX%3A52021PC0206 imposed sanctions and restrictions on the use of such models due to their unexplainable nature.
@@
In recent years, explainable artificial intelligence (XAI) tools have been employed to open the black boxes, in other words, to explain opaque models [7, 8]. These tools have rapidly developed to solve the uninterpretable model structures in finance. The preliminary work by [9] investigated the importance of the variables in the bank failure prediction models, and then [5] obtained the most important variables similarly. [10] explains their financial distress models for the companies to obtain the important variables that can be used to treat loan eligibility requirements. [11] explains the financial distress model to meet the needs of external users generated by local and global explanations. [12] explain the models that are used to predict the financial distress of the companies in China to ensure the reliability of model outputs. While there are studies on explaining black box models in finance, studies on bank failure prediction models have focused only on variable importance. Therefore, it is clear that there is a need for research on explainability in terms of understanding and accountability of such models. Not only have the explainable models, but it is also utilized to derive actionable insights from the models [13, 14]: counterfactual explanations (CE), which describe the changes in the variable values needed to flip the prediction in the intended direction, are used to be actionable. The CE is one of the popular and easy-to-understand XAI tools in many domains such as healthcare [15], education [16, 17], finance [18, 19, 20, 21, 22]. It is an output of an optimization process and is generated with methods such as What-If [23], Nearest instance [24], and Multi-objective counterfactual explanation [25]. The number of CEs that can be generated is potentially limitless [26]; however, generating high-quality CEs can be very difficult [27] because they are quite rare in most cases [28]. The high-quality CEs should be plausible, proximity, sparse, and valid [28]. There are preliminary studies to explore the desired properties of CEs in terms of desiderata. [29] indicated that the quality counterfactuals should be sparse with two variable changes from a psychological point of view. [30] proposed using more plausible CEs instead of proximity CEs because the plausible CEs are more robust. However, no uniformly better CE generation method exists for all datasets from different domains [31]. Thus, it is necessary to evaluate quantitatively [32, 33] the generated CEs by different methods before deployment.
@@
Additionally, the challenges encountered during the model selection phase include poor out-of-time performance and the issue of imbalanced data. The diminished out-of-time performance of bank failure prediction models results in their swift obsolescence [34, 35]. Hence, besides training models with robust out-of-sample performance, ensuring high out-of-time performance for long-term horizon [36, 5, 37] is also imperative. Conversely, the issue of imbalanced data surfaces when there’s a notable scarcity of data examples for failed banks, leading to misclassification by the bank failure prediction model, particularly for the minority class. From the data-centric AI perspective, various resampling methods such as oversampling, undersampling, and Synthetic Minority Over-sampling Technique [38] are employed to mitigate this. However, the efficacy of resampling methods has come under scrutiny in recent years due to their adverse effects on the model [39, 40, 41, 42, 43], sparking debates regarding their utility [44]. Thus, we lean towards using the cost-sensitive approach rather than the resampling methods to address the imbalanced data. Consequently, the impact of resampling techniques on models has been rigorously assessed on both out-of-sample and out-of-time datasets, compared with results obtained using cost-sensitive approaches.
@@
Support vector machines, neural networks, and discriminant analysis are frequently enlisted in bank failure prediction modeling. However, despite their established efficacy in tabular datasets, tree-based models have not garnered widespread adoption, in contrast to deep learning networks [4, 5, 45]. Hence, we offer a novel perspective by leveraging tree-based models, namely decision trees [46], random forests [47], and extra trees [48]. The rationale behind employing multiple models lies in their diverse predictive capabilities, yielding different outcomes with varying variances. This paper solves basic bank failure prediction problems such as imbalance, out-of-sample, and out-of-time performance for US banks. Then, it determines the most appropriate method to generate actionable explanations from black-box bank failure prediction models, considering the effects of the balancing strategies used to produce the highest quality counterfactual explanations. To our knowledge, this is the first paper that proposes using counterfactual explanations to extract actionable insights from the bank failure prediction models. This provides valuable insights to reduce the failure risk of banks, in addition to understanding the model. Moreover, we solve the problems of imbalancedness and out-of-time performance while obtaining the desired counterfactuals. In summary, our contribution is: (1) to train an accurate bank failure prediction model for the US banks against addressing imbalanced data and long time horizon, (2) to obtain the best counterfactual generation method for the bank failure prediction model, and (3) to illustrate how to use the counterfactual explanations to avoid bank failure in practice. This paper is organized as follows. Section 2 provides an overview of the counterfactual explanations, and Section 3 includes information about the conducted experiments on counterfactual generation methods and presents the results. The last section summarises the findings in the paper and suggests further studies.